PLEASE HELP ME I AM NOT GETTING THIS QUESTION....
IF a^2/2=2a,then a equals to
(a)0 or -2
(b)0 or 2
(c)0 or -4
(d)0 or 4
(e)2 or -2 please tell me according to the caps

Answer by Theo(13342)

a^2 / 2 = 2a
multiply both sides of the equation by 2 to get:
a^2 = 4a
divide both sides of the equation by a to get:
a = 4
that's your answer.
a^2 / 2 = 2a becomes:
4^2 / 2 = 2*4 which becomes:
16/2 = 8 which becomes
8=8
answer checks out ok.